OpenStudy (anonymous):

Question 2.2. Find the constant of variation for the relationship shown in the following table: x 1 2 3 4 y 4 8 12 16 (Points : 1) 1 2 3 4

OpenStudy (anonymous):

5. Which does not show a direct variation between x and y? (Points : 1) Y= 4/x y=0.5x y=x/9 y=7x

OpenStudy (anonymous):

help please

OpenStudy (jdoe0001):

x \(\bf {\color{blue}{ 1}}\) 2 3 4 y \(\bf {\color{blue}{ 4}}\) 8 12 16 \(\bf \begin{array}{llll} \textit{something }&\textit{varies directly to }&\textit{something else}\\ \quad \\ \textit{something }&=(\textit{some value })&\textit{something else}\\ \quad \\ y&=(n)&x \end{array}\\ \quad \\ x={\color{blue}{ 1}}\qquad y={\color{blue}{ 4}}\qquad y=(n)x\implies {\color{blue}{ 4}}=(n)\cdot {\color{blue}{ 1}}\) solve for "n", "n" being the "constant of variation"

OpenStudy (anonymous):

so its 4

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

"Which does not show a direct variation between x and y? " well, keep in mind that \(\bf \textit{direct variation}\implies y=(n)x\qquad \textit{inverse variation}\implies y=\cfrac{(n)}{x}\)

OpenStudy (anonymous):

thank u and i no number 2

OpenStudy (jdoe0001):

yw